# Modular knots obey the Chebotarev law

### v3

@@ 6. Modular knots and Lorenz knots
Let ${\mathbb{H}}^{2}=\{z\in{\mathbb{C}}\mid{\rm Im}z>0\}$ denote the upper half plane. The unit tangent bundle of the modular orbifold ${\rm PSL}_{2}{\mathbb{Z}}\setminus\mathbb{H}^{2}$ is well-known to be homeomorphic to both the quotient space ${\rm PSL}_{2}{\mathbb{Z}}\setminus{\rm PSL}_{2}{\mathbb{R}}\cong{\rm SL}_{2}{\mathbb{Z}}\setminus{\rm SL}_{2}{\mathbb{R}}$ and the exterior of a trefoil $K=K_{2,3}$ in $S^{3}$. A flow on ${\rm PSL}_{2}{\mathbb{Z}}\setminus{\rm PSL}_{2}{\mathbb{R}}$ historically called the geodesic flow is defined by multiplying ${\small\begin{pmatrix}e^{t}&0\\ 0&e^{-t}\end{pmatrix}}$ on the right, and its closed orbits are called modular knots around $K_{2,3}$.
@@
The geodesic flow on ${\rm PSL}_{2}{\mathbb{Z}}\setminus{\rm PSL}_{2}{\mathbb{R}}$ is classically known to be Anosov and its structure is finely studied. Note that we have a certain compactification discussed in [BP21]. Unlike the cases of the suspension flows, the flow does not extend to a generalized pseudo-Anosov flow on the result of the 00-filling; in a natural extension of the flow, the missing trefoil consists of four heteroclinic orbits. However, the flow extends to a pseudo-Anosov flow on the result of a generic filling, in which the missing trefoil becomes a closed orbit. Indeed, if we take a compactification or blow-up along the missing trefoil, then we may obtain four attracting meridians on the boundary torus, so it extends to a pseudo-Anosov flow on the $\infty$-surgery.
Thus, by 5 and [McM13, Theorem 1.2], we see the following initial case of 4: modular knots around $K_{2,3}$ and the missing trefoil $K_{2,3}$ in $S^{3}$ obeys the Chebotarev law, if ordered by length in a generic metric.
Here we give some remarks. The classical Lorenz flow in ${\mathbb{R}}^{3}$ [Lor63] is defined by
@@
and its closed orbits in ${\mathbb{R}}^{3}\cup\{\infty\}$ is called Lorenz knots. This flow is well-known for its robustness and the chaotic behavior of its orbits, so that it may suggest the existence of “fate” and “butterfly effects” at the same time. We have a certain geometric model for Lorenz flow as well (cf.[BW83]), and its closed orbits are called Lorenz knots. By virtue of Tucker [Tuc99] and Ghys [Ghy07], the set of modular knots and the missing trefoil in $S^{3}$ is topologically equivalent to that of Lorenz knots, much of whose topological properties having been studied.
The Lorenz flow also has closed orbits that are not Lorenz knots. Conversely, Bonatti–Pinsky [BP21] gives a parametrized family of Lorenz-like flows, and Pinsky [Pin23] points out that some of them contain a sublink that is topologically equivalent to the set of modular knots and the missing trefoil. A similar question to these Lorenz-like knots would be of further interest.
7. The Rademacher symbol
@@
Another distribution theorem is due to Sarnak–Mozzochi, proved by using the trace formula;
Proposition 10 ([Sar08, Sar10, Moz13]).
@@
Exchanging the order of the orbits is an important issue in the context of dynamical systems. The study of Chebyshev bias [RS94] of prime numbers and prime geodesics is of a similar interest. We wonder whether these two equidistribution formulas above may be interpreted from a unified viewpoint.
8. Modular knots around any torus knot
Similar arguments as in Section 6 are applicable to the unit tangent bundles of cusped hyperbolic orbifolds and their covering spaces. As a consequence, we obtain a result on modular knots around any torus knot in $S^{3}$, as follows.
Let $(a,b)$ be the pair of coprime integers and let $\Gamma_{a,b}<{\rm PSL}_{2}{\mathbb{C}}$ denote the $(a,b,\infty)$-triangle group. Then the unit tangent bundle of the orbifold $\Gamma_{a,b}\setminus\mathbb{H}^{2}$ is the exterior of a certain knot $\overline{K}_{a,b}$ in the lens space $L(a,b)$. We have $\pi_{1}L(a,b)={\mathbb{Z}}/r{\mathbb{Z}}$ with $r=ab-a-b$ and the corresponding ${\mathbb{Z}}/r{\mathbb{Z}}$-cover of $L(a,b)-\overline{K}_{a,b}$ is the exterior of the $(a,b)$-torus knot $K_{a,b}$ in $S^{3}$ (cf. [Tsa13]).
There is a standard flow called the geodesic flow in $L(a,b)-\overline{K}_{a,b}$ and its closed orbits are called modular knots in this space. Components of the inverse image of these knots in $S^{3}-K_{a,b}$ are called modular knots around the torus knot $K_{a,b}$.
By a similar argument to [BP21], we see that $\overline{K}_{a,b}$ may consist of several heteroclinic orbits. However, a generic filling of $L(a,b)-\overline{K}_{a,b}$ results in a pseudo-Anosov flow and $\overline{K}_{a,b}$ becomes its closed orbit. So by 5, [McM13, Theorem 1.2], and the latter half of 3, the set of modular knots around $K_{a,b}$ and $K_{a,b}$ obeys the Chebotarev law, as asserted in Theorem 4 in Section 1.
In [MU23], Matsusaka and the author gave a thorough study of modular knots around any torus knot in $S^{3}$, establishing a generalization of Ghys’s formula ${\rm lk}(C_{\gamma},K)=\Psi(\gamma)$, as well as pointing out a version of 10 (see also [vE14, Bur22, BvE22]). This should be compared with Dehornoy–Pinsky’s works on so-called templates and codings [Deh15, DP18]. So, every argument and question in the previous two sections extends to the case around any torus knot.
@@ 9. Further remarks
The $p$-adic torsions of knots and the Lang–Trotter conjecture.
@@
Neukirch–Uchida theorem in anabelian geometry.
In the context of anabelian geometry, the reconstruction problem is one of the main concerns. An important cornerstone is known as Neukirch–Uchida theorem [NSW08, Chapter XII]. In the spirit of arithmetic topology (cf. [Mor12]), we may formulate its analogue for a 3-manifold endowed with a Chebotarev link. Analogues of the Hilbert ramification theory and the Chebotarev density theorem play essential roles there [GUW24].
@@ Acknowledgments
I would like to express my sincere gratitude to Yuta Nozaki, Hirokazu Maruhashi, Kimihiko Motegi, and Toshiki Matsusaka to whom I greatly owe the introductions to low-dimensional generalities, foliated dynamics, non-hyperbolic 0-surgeries, the Rademacher symbols respectively. I also thank Masanori Morishita and Hirofumi Niibo for their cheerful communications during the COVID-19 situation, and Tali Pinsky for answering my query on Lorenz knots. I am grateful to the anonymous referees and experts of the journal for careful reading and sincere comments. The author has been partially supported by JSPS KAKENHI Grant Numbers JP19K14538 and JP23K12969.

### v5

@@ 5. Modular knots and Lorenz knots
We first recall the classical case. Let ${\mathbb{H}}^{2}=\{z\in{\mathbb{C}}\mid{\rm Im}z>0\}$ denote the upper half plane. It is well-known that the unit tangent bundle $T_{1}({\rm PSL}_{2}{\mathbb{Z}}\backslash\mathbb{H}^{2})$ of the modular orbifold ${\rm PSL}_{2}{\mathbb{Z}}\backslash\mathbb{H}^{2}$ is naturally isometric to the quotient space ${\rm PSL}_{2}{\mathbb{Z}}\backslash{\rm PSL}_{2}{\mathbb{R}}\cong{\rm SL}_{2}{\mathbb{Z}}\backslash{\rm SL}_{2}{\mathbb{R}}$ and is homepmorphic to the exterior of a trefoil $K_{2,3}=3_{1}$ in $S^{3}$. The geodesic flow on $T_{1}({\rm PSL}_{2}{\mathbb{Z}}\backslash\mathbb{H}^{2})$ may be seen as the flow on ${\rm SL}_{2}{\mathbb{Z}}\backslash{\rm SL}_{2}{\mathbb{R}}$ defined by multiplying ${\small\begin{pmatrix}e^{t}&0\\ 0&e^{-t}\end{pmatrix}}$ on the right, and its closed orbits are called modular knots around $K_{2,3}$.
@@
The geodesic flow on ${\rm PSL}_{2}{\mathbb{Z}}\backslash{\rm PSL}_{2}{\mathbb{R}}$ is classically known to be Anosov with a dense orbit and its structure is finely studied. We claim that modular knots and the missing trefoil obey the Chebotarev law, if ordered by length in a generic metric.
Unlike the cases of the suspension flows, the geodesic flow does not extend to a generalized pseudo-Anosov flow on the result of the $\infty$-filling resulting in $S^{3}$; in a natural extension of the flow, the missing trefoil consists of four heteroclinic orbits. Even worse, since any surgery along the trefoil in $S^{3}$ results in a non-hyperbolic manifold (cf.[Sul13]), the procedure in the previous sections does not apply. Nevertheless, we may verify this assertion by proving 1.4 stated in Section 1, which is a partial extension of 1.2:
Proof of 1.4.
Let $\Gamma\leq{\rm SL}_{2}{\mathbb{R}}$ be an arbitrary Fuchsian group and let $T_{1}\Sigma$ denote the unit tangent bundle of the orbifold $\Σ\Γ\Sigma=\Gamma\backslash{\mathbb{H}}$, which may have cusps. Note that a lift of the geodesic flow is pseudo-Anosov. Recently, Tsang gave an explicit construction of Markov partition that applies to cusped orbifolds also [Tsa23, Remark 6.7]. By substituting Tsang’s result (see also [KU07] for the case of ${\rm SL}_{2}{\mathbb{Z}}$) to the place in the proof of [McM13, Theorem 1.2] (contained in the proof of [McM13, Theorem 1.4]) where the existence of a Markov partition (a Markov section) for a pseudo-Anosov flow on a closed 3-manifold [FS79] is used, we obtain an extension of 1.2 to such $T_{1}\Sigma$, that is, the closed orbits of the flow obey the Chebotarev law.
(1) If $T_{1}\Sigma=M-L$, then by the definition of the Chebotarev law, we may add the missing finite link $L$ to the family of closed orbits.
(2) If $M-L$ is a finite cover of $T_{1}\Sigma$, then by noting that a Markov partition lifts to the covering space, we obtain a similar result. ∎
As explained in Section 1, for any pair $(a,b)$ of coprime integers, the notion of a modular knot of the $(a,b,\infty)$-triangle group $\Gamma_{a,b}$ is defined.
Proof of 1.5.
1.4 (1) and (2) apply to modular knots in (i) $L(a,b)-\overline{K}_{a,b}$ and (ii) $S^{3}-K_{a,b}$ respectively, yielding the assertion. ∎
Here we give remarks on Lorenz knots. The classical Lorenz flow in ${\mathbb{R}}^{3}$ is defined by
@@
[Lor63]. This flow is well-known for its robustness and the chaotic behavior of its orbits, so it may suggest the existence of “fate” and “butterfly effects” at the same time. We have a certain geometric model for Lorenz flow as well (cf.[BW83]), and its closed orbits are called Lorenz knots. By virtue of Tucker [Tuc99] and Ghys [Ghy07], the set of modular knots and the missing trefoil in $S^{3}$ is topologically equivalent to that of Lorenz knots, much of whose topological properties having been studied.
The Lorenz flow also has closed orbits that are not Lorenz knots. Conversely, Bonatti–Pinsky [BP21] gives a parametrized family of Lorenz-like flows, and Pinsky [Pin23] points out that some of them contain a sublink that is topologically equivalent to the set of modular knots and the missing trefoil. A similar investigation to these Lorenz-like knots would be of further interest.
6. The Rademacher symbols
@@
We remark that Simon recently gave a formula on the linking number of two modular knots [Sim22a, Sim22b], answering Ghys’s question in [Ghy07]. So, we may obtain a similar formula for every fixed modular knot.
Another distribution theorem is due to Sarnak–Mozzochi, proved by using the trace formula:
Theorem 6.2 ([Sar08, Sar10, Moz13]).
@@
Changing the order of the orbits is an important issue in the context of dynamical systems. The study of Chebyshev bias [RS94] of prime numbers and prime geodesics is of a similar interest. We wonder whether these two equidistribution formulas above may be interpreted from a unified viewpoint.
Recently, Matsusaka and the author gave a thorough study of modular knots around any torus knot in $S^{3}$, establishing a generalization of Ghys’s formula ${\rm lk}(C_{\gamma},K)=\Psi(\gamma)$, as well as pointing out a version of 6.2 ([MU23, Theorem A, Corollary 5.3], see also [vE14, Bur22, BvE22]). This should be compared with Dehornoy–Pinsky’s works on so-called templates and codings [Deh15, DP18], as is done by Matsusaka–Shin [MS24]. So, every argument and question for classical modular knots above extends to the case around any torus knot.
@@ 7. Further remarks
Stably Chebotarev
By comparing 1.3 (2) and 1.4 (2), one might wonder if every Chebotarev link is stably Chebotarev, namely, if an infinite link ${\mathcal{L}}$ in a 3-manifold $M$ obeys the Chebotarev law if ordered by length in a generic metric and $h:N\to M$ is a finite branched cover along a finite sublink $L$ of ${\mathcal{L}}$, then the inverse image $h^{-1}({\mathcal{L}})$ has the same property. To discuss this assertion, the Hilbert ramification theory for 3-manifolds ([Uek14], see also [Mor12, Chapter 5]) would play a role, but it turns out to be troublesome, since the covering degrees of the orbits affect the order of the orbits.
The $p$-adic torsions of knots and the Lang–Trotter conjecture
@@
The Neukirch–Uchida theorem in anabelian geometry
In the context of anabelian geometry, the reconstruction problem is one of the main concerns (cf. [Moc15]). An important cornerstone is known as Neukirch–Uchida theorem [NSW08, Chapter XII] for number fields. In the spirit of arithmetic topology, we may formulate its analogue for a 3-manifold endowed with a Chebotarev link. Analogues of the Hilbert ramification theory and the Chebotarev density theorem play essential roles there [GUW24].
@@ Acknowledgments
I would like to express my sincere gratitude to Yuta Nozaki, Hirokazu Maruhashi, Kimihiko Motegi, and Toshiki Matsusaka to whom I greatly owe the introductions to low-dimensional generalities, foliated dynamics, non-hyperbolic 0-surgeries, and the Rademacher symbols respectively. I also thank Masanori Morishita and Hirofumi Niibo for their cheerful communications during the COVID-19 situation, and Tali Pinsky for answering my query on Lorenz knots. I am grateful to the anonymous referees and experts of the journal for careful reading and sincere comments, pointing out several crucial errors in the previous version of this manuscript. The author has been partially supported by JSPS KAKENHI Grant Numbers JP19K14538 and JP23K12969.